# Models, Methods and Waveforms for Estimation and Prediction of Doubly Sparse Time-Varying Channels

### v1

@@ IV-D DS-LTV off-grid channel estimation
where $\mathbf{M}$ and $\mathbf{A}_{\cal P}$ are defined in (16). The minimum mean squared error (MMSE) estimate, $\hat{\boldsymbol{\beta}}$, of $\boldsymbol{\beta}$ based on $\mathbf{y}_{\rm p}$ is given by [27] as
@@
$\sum_{l=0}^{L-1}\frac{1}{N}\mathbb{E}\left[\left\|\mathbf{h}_{l}^{\rm BEM}-\hat{\mathbf{h}}_{l}^{\rm BEM}\right\|^{2}\right]$ $\leq\frac{1}{N}\sum_{l=0}^{L-1}\sum_{q=-Q}^{Q}I_{l,q}\mathbb{E}\left[\left\|\mathbf{h}_{l.q}^{\rm BEM}-\hat{\mathbf{h}}_{l,q}^{\rm BEM}\right\|^{2}\right]$
@@
The value of $N_{\rm p}$ needed for Assumption 3 to hold depends on the particular waveform in use. The numerical results given in Section V show that the relative advantage of AFDM over OFDM and SCM, which has been analytically established by Theorems 1 and 2 under the on-grid channel model, is still valid in the off-grid case.
@@
Under Assumption 3 and Definition 4, provided that $N$ is large enough and $Q_{\rm BEM}\geq C\log\frac{1}{\epsilon}$ for any $\epsilon>0$ and some constant $C$, then $\lim_{\sigma_{w}^{2}\to 0}\mathbb{E}\left[\sum_{l=0}^{L-1}\frac{1}{N}\left\|\mathbf{h}_{l}-\hat{\mathbf{h}}_{l}^{\rm BEM}\right\|^{2}\right]\leq\epsilon$.
@@ IV-E Application to channel extrapolation and prediction
We know from the literature on the Slepian basis [23] that there is a “natural” way to extend the finite sequences $u_{b,n}$ from the smaller interval $\llbracket 0\mkern 1.5mu..\mkern 1.5muN-1\rrbracket$ to the larger one $\llbracket-N_{\text{ext}}\mkern 1.5mu..\mkern 1.5muN+N_{\text{ext}}\rrbracket$, where $N_{\text{ext}}$ denotes the additional channel samples to be extrapolated. This is achieved by letting the index $n$ in (20) be defined over $\mathbb{Z}$ instead of being confined to $\llbracket 0\mkern 1.5mu..\mkern 1.5muN-1\rrbracket$ leading to
@@
Signal $(u_{b,n}^{\rm ext})_{n\in\mathbb{Z}}$ (“$\mathrm{ext}$” stands for “extrapolation”) has a discrete-time Fourier transform (DTFT) that is zero outside $(-W,W)$ and is the signal that has the least energy outside the time interval $\llbracket 0\mkern 1.5mu..\mkern 1.5muN-1\rrbracket$ from among all the discrete-time signals band-limited to $(-W,W)$ [23].
Once we have estimated the multiple-BEM representation of the off-grid DS-LTV channel on the interval $\llbracket 0\mkern 1.5mu..\mkern 1.5muN-1\rrbracket$ as in (31) and once we have calculated the infinite-time version of the DPSS basis function as in (34), the channel can thus be extrapolated as follows
$h_{l,n}^{\rm ext}\triangleq\sum_{q=-Q}^{Q}I_{l,q}\underbrace{e^{\imath 2\pi\frac{nq}{N}}\sum_{b=1}^{Q_{\rm BEM}}\hat{\beta}_{l,q,b}u_{b,n}^{\rm ext}}_{\triangleq h_{l,q,n}^{\rm ext}},\>l=0\cdots L-1\>,\>n\in\mathbb{Z}\>.$ (35)
@@
$h_{l,q,n}^{\rm ext}\triangleq e^{\imath 2\pi\frac{nq}{N}}\sum_{b=1}^{Q_{\rm BEM}}\hat{\beta}_{l,q,b}u_{b,n}^{\rm ext}\>.$ (36)
Defining $\mathbf{u}_{n}^{\rm ext}\triangleq\left[u_{1,n}^{\rm ext}\quad\cdots\quad u_{Q_{\rm BEM},n}^{\rm ext}\right]^{\rm T}$ and referring to (32) gives
$h_{l,q,n}^{\rm ext}=e^{\imath 2\pi\frac{nq}{N}}\left(\mathbf{u}_{n}^{\rm ext}\right)^{\rm T}\mathbf{U}_{Q_{\rm BEM}}^{\rm H}\mathbf{E}_{\frac{q}{N}}^{\rm H}\hat{\mathbf{h}}_{l,q}^{\rm BEM}$ (37)
@@
Therefore, the terms of the sum that contribute the most to the value of $h_{l,q,n}^{\rm ext}=e^{\imath 2\pi\frac{nq}{N}}\sum_{b=1}^{Q_{\rm BEM}}\beta_{l,q,b}u_{b,n}^{\rm ext}$ for $n>N-1$ are the ones with large enough samples $u_{b,n}^{\rm ext}$ i.e., with small enough $\lambda_{b}^{(N,W)}$. While this constrains $Q_{\rm BEM}$ to be sufficiently large, it simultaneously constrains the estimation SNR to be high enough to guarantee a precise estimation of coefficients $\beta_{l,q,b}$ associated with basis functions that have little contribution to the signal received inside the observation interval $\llbracket 0\mkern 1.5mu..\mkern 1.5muN-1\rrbracket$. The issue of setting the value of $Q_{\rm BEM}$ is further discussed in Section V.
The following theorem establishes the optimality in a certain sense of the predictor $h_{l,n}^{\rm ext}$ by relating it to a reduced-rank MMSE estimator.
@@
The predictor $h_{l,n}^{\rm ext}$ defined by (35) for any $n>N-1$ based on extrapolating multiple estimated BEMs of order $Q_{\rm BEM}$ converges in the squared-mean sense in the limit of a vanishing noise variance to a reduced-rank (with a reduced rank equal to $Q_{\rm BEM}$) MMSE estimator of sample $h_{l,n}$ under the assumptions of Definition 4 and Assumption 3.
@@
The DPSS-based BEM modeling in [23] uses a single BEM of order $\check{Q}_{\rm BEM}$ to represent the multi-band signal $h_{l,n}=\sum_{q=-Q}^{Q}I_{l,q}\sum_{i=1}^{N_{\rm D}}\alpha_{l,q,i}e^{\imath 2\pi\frac{n(q+\kappa_{i})}{N}}$ of each channel delay tap to give
@@
where $\check{u}_{b,k}^{(N,W)}$ are the basis vectors of the single multi-band BEM defined as
$\sum_{k=0}^{N-1}\check{C}_{k,n}^{(N,W)}\check{u}_{b,k}=\check{\lambda}_{b}^{(N,W)}\check{u}_{b,n},b=1\cdots N,n=0\cdots N-1\>,$ (39)
@@
The main issue with the multi-band DPSS approach is the size of the codebook that needs to be computed or stored at the network device performing the channel estimation. As can be seen from (29), the codebook size in our approach is the number of columns of the matrix $\mathbf{MB}$ which equals $Q_{\rm BEM}L(2Q+1)$. In the case of the multi-band BEM approach, the codebook size would be $L\sum_{k=1}^{2Q+1}{2Q+1\choose k}\check{Q}_{\rm BEM}(k)\gg Q_{\rm BEM}L(2Q+1)$ as every different combination of $k$ active Doppler grid points would result in a different multi-band prolate matrix $\check{C}_{k,n}^{(N,W)}$ (40) and thus in a different DPSS basis (39). Another advantage of the multiple shifted BEMs scheme is the fact that we can analytically quantify its precision, as we did in Theorem 3 and Corollary 2.

### v2

@@ IV-D DS-LTV off-grid LMMSE channel estimation
where $\mathbf{M}$ and $\mathbf{A}_{\cal P}$ are defined in (19). The MMSE estimate, $\hat{\boldsymbol{\beta}}$, of $\boldsymbol{\beta}$ based on $\mathbf{y}_{\rm p}$ is given by [36] as
@@
Remark 4 (Estimation computational complexity).
The first dominant operation, complexity-wise, in the computation of $\hat{\mathbf{h}}_{l}^{\rm BEM}$ is the matrix inversion in (40). The complexity order of this inversion is cubic in the size of the vector $\boldsymbol{\beta}$, i.e., $\left(s_{\rm d}s_{\rm D}Q_{\rm BEB}\right)^{3}\sim K^{3(\kappa_{\rm d}+\kappa_{\rm D})}$. Here, $s_{\rm d}$ and $s_{\rm D}$ are the delay domain and Doppler domain sparsity levels defined by (11) and (12), respectively, while $\kappa_{\rm d}$ and $\kappa_{\rm D}$ are the parameters of their asymptotic order, as defined by Remark 1. Note that the above value is much smaller than $\left(L(2Q+1)Q_{\rm BEM}\right)^{3}$, the computational complexity of LMMSE channel estimation under no sparsity. The second dominant operation in the computation of $\hat{\mathbf{h}}_{l}^{\rm BEM}$ is matrix $\mathbf{B}$-vector multiplication, an operation that costs $Ns_{\rm d}s_{\rm D}Q_{\rm BEB}\sim K^{2+\kappa_{\rm d}+\kappa_{\rm D}}$. Whether the first or the second operation is the dominant term depends on whether $\kappa_{\rm d}+\kappa_{\rm D}$ is smaller or larger than 1.
@@
$\sum_{l=0}^{L-1}\frac{1}{N}\mathbb{E}\left[\left\|\mathbf{h}_{l}^{\rm BEM}-\hat{\mathbf{h}}_{l}^{\rm BEM}\right\|^{2}\right]$
$\leq\frac{1}{N}\sum_{l=0}^{L-1}\sum_{q=-Q}^{Q}I_{l,q}\mathbb{E}\left[\left\|\mathbf{h}_{l.q}^{\rm BEM}-\hat{\mathbf{h}}_{l,q}^{\rm BEM}\right\|^{2}\right]$
@@
The value of $N_{\rm p}$ needed for Assumption 3 to hold depends on the underlying waveform. The numerical results given in Section V show that the relative advantage of AFDM over OFDM and SCM, which has been analytically established by Theorems 1 and 2 in the on-grid case, is still valid in the off-grid case in the form of better MSE performance at comparable pilot overhead levels.
@@
Under Assumption 3 and Definition 4, provided that $N$ is large enough and $Q_{\rm BEM}\geq C\log\frac{1}{\epsilon}$ for $\epsilon>0$ and some constant $C$, then $\lim_{\sigma_{w}^{2}\to 0}\mathbb{E}[\sum_{l=0}^{L-1}\frac{1}{N}\|\mathbf{h}_{l}-\hat{\mathbf{h}}_{l}^{\rm BEM}\|^{2}]\leq\epsilon$.
@@ IV-E Application to channel extrapolation and prediction
The literature on the Slepian basis [31] establishes a “natural” way to extend the finite sequences $u_{b,n}$ from the smaller interval $\llbracket 0\mkern 1.5mu..\mkern 1.5muN-1\rrbracket$ to the larger one $\llbracket-N_{\text{ext}}\mkern 1.5mu..\mkern 1.5muN+N_{\text{ext}}\rrbracket$, where $N_{\text{ext}}$ denotes the additional channel samples to be extrapolated. This is achieved by letting the index $n$ in (30) be defined over $\mathbb{Z}$ instead of being confined to $\llbracket 0\mkern 1.5mu..\mkern 1.5muN-1\rrbracket$ leading to
@@
Signal $(u_{b,n}^{\rm ext})_{n\in\mathbb{Z}}$ ($\mathrm{ext}$” stands for “extrapolation”) has a discrete-time Fourier transform (DTFT) that is zero outside (W,W)(-W,W) and is the signal that possesses the least energy outside the time interval $\llbracket 0\mkern 1.5mu..\mkern 1.5muN-1\rrbracket$ from among all the discrete-time signals band-limited to (W,W)(-W,W) [31].
Once we have estimated the multiple-BEM representation of the off-grid DS-LTV channel on the interval $\llbracket 0\mkern 1.5mu..\mkern 1.5muN-1\rrbracket$ as in (41) and once we have calculated the infinite-time version of the DPSS basis function as in (44), the channel can thus be extrapolated for $n\in\mathbb{Z}$ as follows
$h_{l,n}^{\rm ext}\triangleq\sum_{q=-Q}^{Q}I_{l,q}\underbrace{e^{\imath 2\pi\frac{nq}{N}}\sum_{b=1}^{Q_{\rm BEM}}\hat{\beta}_{l,q,b}u_{b,n}^{\rm ext}}_{\triangleq h_{l,q,n}^{\rm ext}},\>l=0\cdots L-1\>.$ (45)
@@
$h_{l,q,n}^{\rm ext}\triangleq e^{\imath 2\pi\frac{nq}{N}}\sum_{b=1}^{Q_{\rm BEM}}\hat{\beta}_{l,q,b}u_{b,n}^{\rm ext}\>,n\in\mathbb{Z}\>.$ (46)
Defining $\mathbf{u}_{n}^{\rm ext}\triangleq\left[u_{1,n}^{\rm ext}\quad\cdots\quad u_{Q_{\rm BEM},n}^{\rm ext}\right]^{\rm T}$, (42) gives
$h_{l,q,n}^{\rm ext}=e^{\imath 2\pi\frac{nq}{N}}\left(\mathbf{u}_{n}^{\rm ext}\right)^{\rm T}\mathbf{U}_{Q_{\rm BEM}}^{\rm H}\mathbf{E}_{\frac{q}{N}}^{\rm H}\hat{\mathbf{h}}_{l,q}^{\rm BEM}\>.$ (47)
Remark 6 (Prediction computational complexity).
The dominant cost, in terms of computational complexity, when computing $h_{l,n}^{\rm ext}$ using (45) is the number of complex multiplications needed for the computation of the extrapolated DPSS samples $\{u_{b,n}^{\rm ext}\}_{b=1\cdots Q_{\rm BEM}}$ using (44), which is of the order of $Q_{\rm BEM}N\sim K^{2}$. Note that this complexity order is smaller than that of channel estimation, as established by Remark 4.
@@
Therefore, the terms of the sum that contribute the most to the value of $h_{l,q,n}^{\rm ext}=e^{\imath 2\pi\frac{nq}{N}}\sum_{b=1}^{Q_{\rm BEM}}\beta_{l,q,b}u_{b,n}^{\rm ext}$ for n>N1n>N-1 are the ones with large enough $u_{b,n}^{\rm ext}$, i.e., with small enough $\lambda_{b}^{(N,W)}$. While this constrains $Q_{\rm BEM}$ to be sufficiently large, it also constrains the SNR to be high enough to guarantee a precise estimation of the coefficients $\beta_{l,q,b}$ associated with basis functions that have little contribution to the signal received within the observation interval $\llbracket 0\mkern 1.5mu..\mkern 1.5muN-1\rrbracket$. The issue of setting the value of $Q_{\rm BEM}$ is further discussed in Section V. Theorem 4 establishes the optimality of the predictor $h_{l,n}^{\rm ext}$ by relating it to reduced-rank MMSE.
@@
Under the assumptions of Definition 4 and Assumption 3 and for any n>N1n>N-1, the predictor $h_{l,n}^{\rm ext}$ defined by (45) converges, in the squared-mean sense in the limit of a vanishing noise variance, to a reduced-rank MMSE estimator of $h_{l,n}$ with a reduced rank equal to $Q_{\rm BEM}$.
@@
The DPSS-based multi-band single-BEM modeling in [31] uses a single BEM of order $\check{Q}_{\rm BEM}$ to represent the multi-band signal $h_{l,n}=\sum_{q=-Q}^{Q}I_{l,q}\sum_{i=1}^{N_{\rm D}}\alpha_{l,q,i}e^{\imath 2\pi\frac{n(q+\kappa_{i})}{N}}$ of each channel delay tap to give
@@
where $\check{u}_{b,k}^{(N,W)}$ are the basis vectors of the single multi-band BEM defined for $b=1,\ldots,N$ as
$\sum_{k=0}^{N-1}\check{C}_{k,n}^{(N,W)}\check{u}_{b,k}=\check{\lambda}_{b}^{(N,W)}\check{u}_{b,n},n=0\cdots N-1\>,$ (49)
@@
In terms of computational complexity, the extrapolation of the multi-band single-BEM approach as done in (51) is equivalent to the extrapolation step of the “multiple shifted elementary BEMs” approach in (45). The main issue with the multi-band single-BEM approach is the size of the codebook that needs to be computed or stored at the network device performing channel estimation. As can be seen from (39), the codebook size in our approach is the number of columns of the matrix $\mathbf{MB}$ which equals $Q_{\rm BEM}L(2Q+1)$. In the case of the multi-band single-BEM approach, the codebook size is $L\sum_{k=1}^{2Q+1}{2Q+1\choose k}\check{Q}_{\rm BEM}(k)\gg Q_{\rm BEM}L(2Q+1)$ as every different combination of $k$ active Doppler grid points would result in a different multi-band prolate matrix $\check{C}_{k,n}^{(N,W)}$ (50) and thus in a different DPSS basis (49).